Systems Staff Meeting Notes
July 30, 2007
July 30, 2007
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Bill Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Judith Russell, Mark Sullivan, Stephen WilliamsAnnouncements
Ongoing Systems Updates
- Amy's dog learned to speak on command.
- Question & answer session with Dean Judy Russell took place at the conclusion of this staff meeting.
- Orders - nothing new.
- Supply budget - Will still needs to do the training before the end of August so he will have his own PCard.
- Training is on the calendar for August 20 - may not be a good date as it is the first week of the fall semester.
- Bobbie notes that she still has not received updates from the liaisons as to where the new machines have been set up. Amy will ask liaisons to send Bobbie an updated spreadsheet.
Software Projects Status
- Exchange server - Finished working on it this morning. There were a lot of log errors. Exchange could not contact the domain controller. Will called Dell. For some reason, it thought the IPs were reversed for DC1 and DC2. Exchange is set for the third setting - Campus DNS. It was looking for DC1 at the DC2 address. Conflicts should be gone now. The Dell DNS expert called this morning. The management tool works now. The old, incorrect records were deleted. Will is going to test it tonight and fail it over.
- Mobile device synch and Web mail problems - there was a misconfiguration in the IIS settings. Will thinks it's fixed.
- NT4 - problem over the weekend - it dropped off the domain and broke Verify Patron. The failover worked in most places, but some with a new image had the ini in the wrong place. The final test on the new image had not taken place. The launch from the registry has to be in one location. NT4 is now back on the domain. We had difficulty finding the admin password - had to purchase a tool to retrieve it.
- Public computer replacements - most have been done, only the smaller departments remain: Special Collections, Education, LAC.
- Xerox - nothing new. Format changes in PDFs will cause printing problems.
- SAN is racked. Logan will attend VM Ware training next week, Tues-Fri.
- SMS - Logan talked to CNS about DHCP settings for remote deployment. CNS will need to make changes to DHCP in West.
- Server updates - okay.
- Grover design review - Chris and Stephen are working on recommended changes.
- DLC Tracking - making changes: a lot of tools - interface - importers. Different countries/departments - we need to track where the changes originate.
- Greenstone - hardware - Logan is analyzing the possible use of the excess storage.
- Greenstone - Mark is building the billing authority database and working on general workflow materials and archive.
- Conference call with Atlas - Lori Driscoll, Winston and Michelle Foss - the switch from barcode to UFID is scheduled for August 10th. Winston sent a crosswalk table about two months ago, will test it after the switch.
- Patron Updates - encrypted files sent to FCLA. Friday - the correct file was not loaded by FCLA. We will send the files again on Thursday. We will do it again after the version 18 change.
Students and Department
- Two or three weeks ago, we put up a survey regarding the food policy in the Library. We are getting a lot of feedback.
- Barbara Hood is talking about redoing the fundraising site. She has a volunteer to do it for her. She is currently gathering ideas. In the fall semester, it will be subject to the UF Web page guidelines.
- Capital campaign upcoming.
- In 2-3 weeks, Endeca will be coming out. All Catalog links will have to be changed. Global changes should work. Will investigate doing it from the service side.
- Homepage design - need new template.
- UF Web committee - document on new identity - new header would be best.
- Bill had some suggestions for search on the Homepage.
Systems Security Management
- Vera is gone.
- Mindy will start coming in at 11AM.
- Luiza will return in the fall.
- Meeting held.